Suppose a group of 800 smokers (who all wanted to give up smoking) were randomly assigned to receive an antidepressant drug or a placebo for six weeks. Of the 116 patients who received the antidepressant drug, 21 were not smoking one year later. Of the 684 patients who received the placebo, 39 were not smoking one year later. Given the null hypothesis H0:(pdrug−pplacebo)=0H0:(pdrug−pplacebo)=0 and the alternative hypothesis Ha:(pdrug−pplacebo)≠0Ha:(pdrug−pplacebo)≠0, conduct a test to see if taking an antidepressant drug can help smokers stop smoking. Use α=0.04α=0.04,

(a) The test statistic is

(b) The P-value is

(c) The final conclusion is

A. There is not sufficient evidence to determine whether the antidepressant drug had an effect on changing smoking habits after one year.
B. There seems to be evidence that the patients taking the antidepressant drug have a different success rate of not smoking after one year than the placebo group.
Answer #1

a)


DrugPlacebo
x=   2139
p̂=x/n=0.18100.0570
n =116684
estimated prop. diff =p̂1-p̂2    =0.1240
pooled prop p̂ =(x1+x2)/(n1+n2)=0.0750



std error Se=√(p̂1*(1-p̂1)*(1/n1+1/n2) =0.0264
test stat z=(p̂1-p̂2)/Se =4.69

b)

P value   =0.0000(from excel:2*normsdist(-4.69)

c)

since p value <0.04:

B. There seems to be evidence that the patients taking the antidepressant drug have a different success rate of not smoking after one year than the placebo group.


answered by: ANURANJAN SARSAM
Answer #2

a)

Drug Placebo
x=    21 39
p̂=x/n= 0.1810 0.0570
n = 116 684
estimated prop. diff =p̂1-p̂2    = 0.1240
pooled prop p̂ =(x1+x2)/(n1+n2)= 0.0750
std error Se=√(p̂1*(1-p̂1)*(1/n1+1/n2) = 0.0264
test stat z=(p̂1-p̂2)/Se = 4.69

b)

P value   = 0.0000 (from excel:2*normsdist(-4.69)

c)

since p value <0.04:

B. There seems to be evidence that the patients taking the antidepressant drug have a different success rate of not smoking after one year than the placebo group.